Grade 12Algebra

Check whether the value given in the brackets is a solution to the given equation or not:
n + 5 = 19 (n = 1)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

LHS = n + 5
By substituting the value of n = 1
Then,
LHS = n + 5
= 1 +5
= 6
By comparing LHS and RHS
6≠19
LHS≠RHS
Hence, the value of n = 1 is not a solution to the given equation n + 5 = 19.

​Thanks